Find the equation of a line passing through (3,6) and parallel to the line y+x=7
Answer by stanbon(75887)

Find the equation of a line passing through (3,6) and parallel to the line y+x=7
----------------
"Parallel to the line y=-x+7"
That line has slpe = -1
---------
Equation has slope=-1,x=3,y=6
y=mx+b
6=-1*3+b
b=9
EQUATION:
y=-x+9
Cheers,
Stan H.